OpenStudy (anonymous):

MEDALS GIVEN Suppose your business has a special checking account used just for paying the phone bill. The balance is $740.00 this month. Next month the balance will be $707.60, after that it will be $675.20, and on the third month the balance will be $642.80. Write a rule to represent the balance in the account as an arithmetic sequence. How many months can you pay your phone bill without depositing any more money in the account?

OpenStudy (anonymous):

A(n)=740.00+(n-1)(-32.40);23months A(n)=740.00-32.40n,23 months A(n)=740.00+(n+1)(-32.40);24 months